You wish to test the following claim (HaHa) at a significance level of α=0.005.

      Ho:p1=p2
      Ha:p1>p2

You obtain 106 successes in a sample of size n1=522 from the first population. You obtain 33 successes in a sample of size n2=270 from the second population. For this test, you should NOT use the continuity correction, and you should use the normal distribution as an approximation for the binomial distribution.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic = _______________

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value = ____________________

The p-value is...

  • ____________less than (or equal to) αα
  • ____________greater than αα



This test statistic leads to a decision to...

  • ______reject the null
  • ________accept the null
  • ________fail to reject the null



As such, the final conclusion is that...

  • _________There is sufficient evidence to warrant rejection of the claim that the first population proportion is greater than the second population proportion.
  • _________There is not sufficient evidence to warrant rejection of the claim that the first population proportion is greater than the second population proportion.
  • _________The sample data support the claim that the first population proportion is greater than the second population proportion.
  • _________There is not sufficient sample evidence to support the claim that the first population proportion is greater than the second population proportion.
Answer #1

Null and alternative hypothesis: H. :P = P2 H :P > P2 Significance level: a =0.005 The sample proportions are În = 1 - 100 =Test statistic: - r- i 1 17(1-7) +1 Inn 0.2031-0.1222 10.1755(1–0.1755) 5221 ) = 2.835 P-value: p= P(Z >=) =1-P(Z < 2.835) =1Decision rule: Reject H, if psa otherwise do not reject H. The p-value is less than the level of significance. This test stat